# Cross-Modal Taxonomic Generalization in (Vision-) Language Models

### 08 Mar 2026 v1

@@ 2 Related Work
Our findings also relate to recent proposals about the convergence in the representations of models trained on different modalities or domains (Cao and Yamins, 2021; Sorscher et al., 2022; Huh et al., 2024; Wu et al., 2025). For instance, according to the ‘Platonic Representation Hypothesis’ (Huh et al., 2024), the representations learned by unimodal models of different modalities are converging since they are statistical models of the same world. We shed light on the implications of these proposals when two unimodal models (the vision encoder and the LM) are aligned. When this alignment happens under counterfactual scenarios (Section˜4.3), where the incoming input signals are at odds with the conceptual organization of the two models—posited to be similar under the aforementioned proposals—we see a severe drop in taxonomic generalization. This underscores the importance of such representational convergence in the extreme setting where explicit information about the phenomenon (here, hypernymy) is restricted to a model from one modality.
@@ 4.2 Cross-Modal Taxonomic Generalization
Having shown that the LM backbones capture hypernymy, that the trained VLMs can generalize to unseen images regardless of whether or not the image encoder receives textual supervision, and that they have learned to perform the task, we now turn to our main result, shown in the rightmost column of Figure˜4, named “Held-Out Hypernyms”. We see that regardless of the amount of exposure to a hypernym category, LM backbones with pre-trained representations achieve consistently above-chance performance in predicting the presence of hypernyms in images not seen during the models’ training. This is observed for both types of ablations, and even at 100% ablation, where the model never sees any hypernym label during training. Models with randomly initialized LM backbone representations consistently underperform their pre-trained counterparts, and are close to chance in the absence of any hypernym-level supervision. These results also hold for Llama 3.2, which we report in Figure˜10 in the Appendix. These results suggest that the pre-trained representations of LMs can facilitate cross-modal taxonomic generalization.

### 22 Apr 2026 v2

@@ 2 Related Work
Our findings also relate to recent proposals about the convergence in the representations of models trained on different modalities or domains (Cao and Yamins, 2021; Sorscher et al., 2022; Li et al., 2024; Huh et al., 2024; Wu et al., 2025). For instance, according to the ‘Platonic Representation Hypothesis’ (Huh et al., 2024), the representations learned by unimodal models of different modalities are converging since they are statistical models of the same world. We shed light on the implications of these proposals when two unimodal models (the vision encoder and the LM) are aligned. When this alignment happens under counterfactual scenarios (Section˜4.3), where the incoming input signals are at odds with the conceptual organization of the two models—posited to be similar under the aforementioned proposals—we see a severe drop in taxonomic generalization. This underscores the importance of such representational convergence in the extreme setting where explicit information about the phenomenon (here, hypernymy) is restricted to a model from one modality.
@@ 4.2 Cross-Modal Taxonomic Generalization
Having shown that the LM backbones capture hypernymy, that the trained VLMs can generalize to unseen images regardless of whether or not the image encoder receives textual supervision, and that they have learned to perform the task, we now turn to our main result, shown in the rightmost column of Figure˜4, named “Held-Out Hypernyms”. We see that regardless of the amount of exposure to a hypernym category, LM backbones with pre-trained representations achieve consistently above-chance performance in predicting the presence of hypernyms in images not seen during the models’ training. This is observed for both types of ablations, and even at 100% ablation, where the model never sees any hypernym label during training. Models with randomly initialized LM backbone representations consistently underperform their pre-trained counterparts, and are close to chance in the absence of any hypernym-level supervision. These results also hold for Llama 3.2, which we report in Figure˜10 in the Appendix. These results suggest that the pre-trained representations of LMs can facilitate cross-modal taxonomic generalization. We analyze how this generalization varies across depths of taxonomic organization in Appendix˜F.
@@ Appendix E Results on Qwen3-8B LM backbone
Appendix F Cross-Modal Generalization at Varying Depths of Taxonomic Organization
Our set of 53 categories can be organized hierarchically into multiple tree-like structures, each representing a taxonomy of a specific higher level category. The maximum depth across these trees is 3 (e.g., kitchen tool $\prec$ tool $\prec$ hardware, where $\prec$’ represents the hypernym relation, with the right hand side category being the hypernym of the left hand side category). Overall, we have 41 categories at depth 1, 9 categories at depth 2, and 3 categories at depth 3. To understand how cross-modal taxonomic generalization of our tested models varies across these depths, we present results (here, at the 100% ablation level) divided across the three depths for all models in Figure˜12. We see that models generally perform better at depth 1 and 2 than at depth 3, i.e., they are markedly better at generalizing to lower level categories than at higher level categories.
Note that it is not clear if the poorer performance at depth 3 is necessarily because the models struggle at multi-hop reasoning (over hierarchically organized category structures), since the internal mechanism of how models arrive at their answer is not obvious. An alternate explanation could be the ‘frequency of instantiation’ hypothesis (Barsalou, 1985), which states that the number of times a learner observes a given category as member of another category determines the strength of the category relationship between the two. Therefore, the LM backbones may have encountered fewer instances of hyponym/leaf concepts as members of depth 3 hypernyms relative to depth 1 and 2 hypernyms. This does not appeal directly to a process that commits to an explicit hierarchical representation of categories, and yet can explain the observed result. We therefore caution readers from making inferences about the nature of explicit multi-hop hierarchical reasoning in the models from these results.
Figure 12: Macro F1s on unseen images for all four LM backbones across various depths of category organization in our analysis, computed for the 100% ablation experiment. Error bars indicate 95% confidence intervals across three random seeds and number of categories per depth ($N_{1}$ = 41; $N_{2}$ = 9; $N_{3}$ = 3).